# RepLoRA: Reparameterizing Low-Rank Adaptation via the Perspective of Mixture of Experts

### v1

@@ 7 Conclusion
We introduced a theoretical framework that bridges LoRA with MoE, offering new insights into the benefits of reparameterizing LoRA for achieving optimal sampling efficiency. Building on this foundation, we proposed RepLoRA, an effective and efficient PEFT approach. To evaluate RepLoRA, we conducted extensive experiments across four domains: image, video, text, and multimodal tasks. RepLoRA substantially outperformed LoRA and other PEFT methods in all settings, demonstrating its adaptability and effectiveness. These results highlight the potential of reparameterized structures in enhancing efficiency and effectiveness for fine-tuning large-scale models.
Supplement to “RepLoRA: Reparameterizing Low-Rank Adaptation via the Perspective of Mixture of Experts”
In this supplementary material, we provide proofs of the main results in Appendix A. Related works on parameter-efficient fine-tuning techniques, low-rank adaptation, and the mixture of experts are discussed in Appendix B. Details of experiments are in Appendix C while additional experiments are in Appendix D.
Appendix A Proofs
@@
The proof of Theorem 4.1 consists of two key steps.
Step 1. In this step, we demonstrate that the following limit holds for any $r\geq 1$:
@@
To prove the inequality above, it is sufficient to construct a sequence of mixing measures $(G_{n})_{n\geq 1}$ satisfying
$\lim_{n\to\infty}\mathcal{D}_{1,r}(G_{n},G_{*})=0$ and $\lim_{n\to\infty}{\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}}/{\mathcal{D}_{1,r}(G_{n},G_{*})}=0.$
Indeed, we consider the following sequence $G_{n}=\sum_{i=1}^{L+1}\exp(c^{n}_{i})\delta_{({\bm{B}}^{n}_{Q,i},{\bm{A}}^{n}_{Q,i},{\bm{B}}^{n}_{V,i},{\bm{A}}^{n}_{V,i})}$, where
@@
in which without loss of generality, we assume that $({\bm{A}}^{*}_{V,1})^{(1)}\neq 0$. With this definition, the loss function $\mathcal{D}_{1,r}(G_{n},G_{*})$ can be computed as
@@
Therefore, $\mathcal{D}_{1,r}(G_{n},G_{*})\to 0$ as $n\to\infty$.
We will show that $\lim_{n\to\infty}\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{1,r}(G_{n},G_{*})=0$. Consider the quantity
@@
$=\frac{1}{2}\Big{[}\exp(b_{*,1})+\frac{1}{n^{r+1}}\Big{]}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{B}}_{Q,1}^{*}{\bm{A}}_{Q,1}^{*})\mathbb{X})[({\bm{B}}^{n}_{V,1}{\bm{A}}^{n}_{V,1}-{\bm{B}}^{*}_{V,1}{\bm{A}}^{*}_{V,1})$
$+({\bm{B}}^{n}_{V,2}{\bm{A}}^{n}_{V,2}-{\bm{B}}^{*}_{V,1}{\bm{A}}^{*}_{V,1})]\mathbb{X}$
@@
Moreover, we can also verify that $B_{n}(\mathbb{X})=0$, and $C_{n}(\mathbb{X})=\mathcal{O}(n^{-(r+1)})$. Therefore, for almost every $\mathbb{X}$, it can be deduced that
$\lim_{n\to\infty}Q_{n}(\mathbb{X})/\mathcal{D}_{1,r}(G_{n},G_{*})=0.$
Notice that the term $\Big{[}\sum_{k=1}^{L}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{B}}_{Q,k}^{*}{\bm{A}}_{Q,k}^{*})\mathbb{X}+c^{*}_{k})\Big{]}$ is bounded. Thus, for almost every $\mathbb{X}$, we have
$\frac{f_{G_{n}}(\mathbb{X})-f_{G_{*}}(\mathbb{X})}{\mathcal{D}_{1,r}(G_{n},G_{*})}\to 0.$
This limit suggests that
$\lim_{n\to\infty}\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{1,r}(G_{n},G_{*})=0.$
Therefore, we obtain the claim in equation (19).
Step 2. Next, we will demonstrate that
@@
From the hypothesis, $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{n}$ are i.i.d. from a Gaussian distribution. Therefore, we obtain that the conditional distribution of $Y_{i}|\mathbb{X}_{i}$ is also a Gaussian distribution, namely, we have $Y_{i}|\mathbb{X}_{i}\sim\mathcal{N}(f_{G_{*}}(\mathbb{X}_{i}),\sigma^{2})$ for all $i\in[n]$. Furthermore, when $\varepsilon>0$ is sufficiently small, given a fixed constant $C_{1}>0$ (the specific value of $C_{1}$ will be chosen later) there exists $G^{\prime}_{*}\in\mathcal{G}_{L^{\prime}}(\Theta)$ such that the following results hold: $\mathcal{D}_{1,r}(G^{\prime}_{*},G_{*})=2\varepsilon$ and $\|f_{G^{\prime}_{*}}-f_{G_{*}}\|_{L^{2}(\mu)}\leq C_{1}\varepsilon$, which is due to the result in equation (19). An application of the Le Cam’s lemma [54] along with the weak triangle inequality property of the Voronoi loss function $\mathcal{D}_{1,r}$ leads to
@@
$\gtrsim\varepsilon\cdot\text{exp}(-C_{1}n\varepsilon^{2}).$ (22)
In these inequalities, we utilize the following well-known closed-form expression for the KL divergence between two Gaussian distributions:
@@
By choosing $\varepsilon=n^{-1/2}$, we can verify that $\varepsilon\cdot\text{exp}(-C_{1}n\varepsilon^{2})=n^{-1/2}\exp(-C_{1})$. As a consequence, the minimax lower bound in equation (21) is achieved and the proof is completed.
@@
We start with the following position, which establishes the convergence rate of the regression function estimation $f_{\bar{G}_{n}}$ to the true regression function $f_{\bar{G}_{*}}$:
@@
The proof of that inequality will consist of two parts:
  • Local part: $\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,\bar{G}_{*})\leq\varepsilon}\|f_{G}-f_{\bar{G}_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{2}(G,\bar{G}_{*})>0.$
  • Global part: $\inf_{G\in\bar{\mathcal{G}}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,\bar{G}_{*})\leq\varepsilon^{\prime}}\|f_{G}-f_{\bar{G}_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{2}(G,\bar{G}_{*})>0.$
@@ A.2.1 Local Part
We first consider the proof for the local part. We will prove that
@@
Assume, on the contrary, that the claim does not hold. It indicates that we can find a sequence of mixing measures $G_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(c_{n,j^{\prime}})\delta_{{\bm{B}}_{n,j^{\prime}}{\bm{A}}_{n,j^{\prime}}}$ in $\bar{\mathcal{G}}_{L^{\prime}}(\Theta)$ such that
@@
Since $\mathcal{D}_{2n}\to 0$, we have $\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\to\exp(c_{*,j})$, ${\bm{Z}}_{n,i}\to{\bm{Z}}_{j}^{*}$ for any index $i$ in the Voronoi cell $\mathcal{V}_{j}$ and for any index $j\in[L]$. Throughout this proof, we assume without loss of generality that $M^{0}_{K,j}=I_{\bar{d}}$ with a note that our techniques can be extended to the general setting of that matrix. Now, the proof of the local part is divided into three steps as follows:
@@
Then, via simple algebra the function $Q_{n}(\mathbb{X})$ can be decomposed into three terms as follows:
@@
We denote $\bar{U}(\mathbb{X};{\bm{Z}}):=\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}})\mathbb{X})$ and $\bar{V}(\mathbb{X};{\bm{Z}})=({\bm{M}}^{0}_{V}+{\bm{Z}})\mathbb{X}$, and $F(\mathbb{X};{\bm{Z}})=\bar{U}(\mathbb{X};{\bm{Z}})\bar{V}(\mathbb{X};{\bm{Z}})$. Our strategy is to decompose the function $\bar{A}_{n}(\mathbb{X})$ into two parts: the first part consists of Voronoi cells with only one element and the second part comprises of Voronoi cells with more than one element. In summary, the function $\bar{A}_{n}(\mathbb{X})$ has the following decomposition:
@@
We aim to utilize first-order Taylor expansion to the function $F$ in the formulation of the function $\bar{A}_{n,1}(\mathbb{X})$. It can be done by employing the first-order Taylor expansions to each function $\bar{U}$ and $\bar{V}$, which are given by:
@@
for any $i$ and $j$ such that $i\in\mathcal{V}_{j}$ and $|\mathcal{V}_{j}|=1$. Here, both the functions $\bar{R}_{ij,1}(\mathbb{X})$ and $\bar{R}_{ij,2}(\mathbb{X})$ in the Taylor expansions of the function $\bar{U}$ and $\bar{V}$ denote the Taylor remainders. Combining all the above results together, we arrive at
@@
where the function $\bar{R}_{n,1}(\mathbb{X})$ satisfies that $\bar{R}_{n,1}(\mathbb{X})/\mathcal{D}_{2n}\to 0$. It is due to the uniform Lipschitz property of the function $F$. In the formulation of the function $\bar{A}_{n,1}(\mathbb{X})$, the coefficients $\bar{M}_{n,j,\alpha}$ admit the following forms:
@@
Moving to the function $\bar{A}_{n,2}(\mathbb{X})$, we utilize the Taylor expansion to decompose that function. However, as each Voronoi cell in the formulation of the function $\bar{A}_{n,2}(\mathbb{X})$ has more than one element, we need to use second-order Taylor expansion to account for the convergence of all the elements in those Voronoi cells to the same limit point. Given that idea, we apply the second-order Taylor expansion and achieve that
@@
$\bar{A}_{n,1}(\mathbb{X})=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{V}+{\bm{Z}}_{j}^{*})\mathbb{X}$
$\hskip 56.9055pt+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}\Big{]}+\bar{R}_{n,1}(\mathbb{X}),$
$\bar{A}_{n,2}(\mathbb{X})=\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{V}+{\bm{Z}}_{j}^{*})\mathbb{X}$
$\hskip 56.9055pt+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}$
@@
$Q_{n}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})$ (25)
$+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X}$ (26)
$+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}$
@@
Now, for any indices $j\in[L]$ such that $|\mathcal{V}_{j}|=1$, the limits $\bar{M}_{n,j,e_{uv}}/\mathcal{D}_{2n}\to 0$ lead to
$\frac{\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\|\Delta{\bm{Z}}_{n,ij}\|_{1}}{\mathcal{D}_{2n}}\to 0.$
That result directly implies that
@@
By putting the results in equations (28), (29), and (30) together, we arrive at $1=\frac{\mathcal{D}_{2n}}{\mathcal{D}_{2n}}\to 0$ as $n\to\infty$, which is a contradiction. Consequently, as $n\to\infty$, at least one of the coefficients of the linear independent functions in $Q_{n}(\mathbb{X})/\mathcal{D}_{2n}$ does not go to 0 .
Step 3 - Applying the Fatou’s lemma.
Denote $\bar{m}_{n}$ the maximum of the absolute values of the coefficients of the linear independent functions in $Q_{n}(\mathbb{X})/\mathcal{D}_{2n}$. As at least one of these coefficients does not go to 0, it indicates that $1/\bar{m}_{n}\not\to\infty$ as $n\to\infty$. Since $\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{2n}\to 0$ as $n\to\infty$, we obtain $\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/(\bar{m}_{n}\mathcal{D}_{2n})\to 0$. An important insight is that we can move the limit as $n\to\infty$ of that ratio inside the integral via an application of the Fatou’s lemma, which is given by:
[…]
$\geq\int\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}(\mathbb{X})-f_{\bar{G}_{*}}(\mathbb{X})\right|}{\bar{m}_{n}\mathcal{D}_{2n}}d\mu(\mathbb{X})\geq 0.$

### v2

@@ 8 Conclusion
We introduced a theoretical framework that bridges LoRA with MoE, offering new insights into the benefits of reparameterizing LoRA for achieving optimal sampling efficiency. Building on this theoretical foundation, we proposed RepLoRA, an effective and efficient approach to PEFT. To evaluate RepLoRA, we conducted extensive experiments across four diverse domains: image, video, text, and multimodal tasks. RepLoRA substantially outperformed LoRA and other PEFT methods in all settings, demonstrating its adaptability and effectiveness. These results highlight the potential of reparameterized structures in enhancing efficiency and effectiveness for fine-tuning large-scale models.
Impact Statement
This paper presents work that aims to advance the field of Machine Learning. Our work has potential societal consequences, none of which we feel must be specifically highlighted here.
Acknowledgement
Trung Le was partly supported by ARC DP23 grant DP230101176 and by the Air Force Office of Scientific Research under award number FA9550-23-S-0001.
Appendix A Proofs of Theoretical Results
@@
The proof is divided into two steps as follows:
Step 1. To begin with, we demonstrate that the following limit holds for any $r\geq 1$:
@@
Note that it is sufficient to construct a mixing measure sequence $(G_{n})_{n\geq 1}$ that satisfies both $\mathcal{D}_{1,r}(G_{n},G_{*})\to 0$ and ${\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}}/{\mathcal{D}_{1,r}(G_{n},G_{*})}\to 0,$ as $n\to\infty$.
For that purpose, we take into account the sequence $G_{n}=\sum_{i=1}^{L+1}\exp(c^{n}_{i})\delta_{({\bm{B}}^{n}_{Q,i},{\bm{A}}^{n}_{Q,i},{\bm{B}}^{n}_{V,i},{\bm{A}}^{n}_{V,i})}$, where
@@
in which we assume WLOG that $({\bm{A}}^{*}_{V,1})^{(1)}\neq 0$.
Then, we can compute the loss function $\mathcal{D}_{1,r}(G_{n},G_{*})$ as
@@
It can be seen that $\mathcal{D}_{1,r}(G_{n},G_{*})\to 0$ as $n\to\infty$.
Subsequently, we illustrate that $\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{1,r}(G_{n},G_{*})\to 0$. In particular, let us consider the quantity
@@
$=\frac{1}{2}\Big{[}\exp(b_{*,1})+\frac{1}{n^{r+1}}\Big{]}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{B}}_{Q,1}^{*}{\bm{A}}_{Q,1}^{*})\mathbb{X})[({\bm{B}}^{n}_{V,1}{\bm{A}}^{n}_{V,1}-{\bm{B}}^{*}_{V,1}{\bm{A}}^{*}_{V,1})+({\bm{B}}^{n}_{V,2}{\bm{A}}^{n}_{V,2}-{\bm{B}}^{*}_{V,1}{\bm{A}}^{*}_{V,1})]\mathbb{X}$
@@
Moreover, we can also verify that $B_{n}(\mathbb{X})=0$, and $C_{n}(\mathbb{X})=\mathcal{O}(n^{-(r+1)})$. Thus, we deduce that $Q_{n}(\mathbb{X})/\mathcal{D}_{1,r}(G_{n},G_{*})\to 0$ as $n\to\infty$ for almost every $\mathbb{X}$.
As the term $\Big{[}\sum_{k=1}^{L}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{B}}_{Q,k}^{*}{\bm{A}}_{Q,k}^{*})\mathbb{X}+c^{*}_{k})\Big{]}$ is bounded, we have $[f_{G_{n}}(\mathbb{X})-f_{G_{*}}(\mathbb{X})]/\mathcal{D}_{1,r}(G_{n},G_{*})\to 0$ for almost every $\mathbb{X}$. This limit suggests that
$\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{1,r}(G_{n},G_{*})\to 0$
as $n\to\infty$. Thus, we obtain the claim in equation (19).
Step 2. We will establish the desired result in this step, that is,
@@
Since the noise variables $\epsilon_{i}$ follow from the Gaussian distribution, we get that $Y_{i}|\mathbb{X}_{i}\sim\mathcal{N}(f_{G_{*}}(\mathbb{X}_{i}),\sigma^{2})$ for all $i\in[n]$. Additionally, for sufficiently small $\varepsilon>0$ and a fixed constant $C_{1}>0$ which we will select later, we can find a mixing measure $G^{\prime}_{*}\in\mathcal{G}_{L^{\prime}}(\Theta)$ such that $\mathcal{D}_{1,r}(G^{\prime}_{*},G_{*})=2\varepsilon$ and $\|f_{G^{\prime}_{*}}-f_{G_{*}}\|_{L^{2}(\mu)}\leq C_{1}\varepsilon$ thanks to the result in equation (19). According to the Le Cam’s lemma (Yu, 1997), as the Voronoi loss function $\mathcal{D}_{1,r}$ satisfies the weak triangle inequality, it follows that
@@
$\gtrsim\varepsilon\cdot\text{exp}(-C_{1}n\varepsilon^{2}),$ (22)
where the second inequality follows from the equality
@@
Let $\varepsilon=n^{-1/2}$, then we get that $\varepsilon\cdot\text{exp}(-C_{1}n\varepsilon^{2})=n^{-1/2}\exp(-C_{1})$. Consequently, we achieve the desired minimax lower bound in equation (21). Related works on parameter-efficient fine-tuning techniques, low-rank adaptation, and mixture of experts are in Appendix
@@
We first start with the following result regarding the convergence rate of the regression function estimation $f_{\bar{G}_{n}}$ to the true regression function $f_{\bar{G}_{*}}$:
@@
We divide the proof of the above inequality into local and global parts.
@@ A.2.1 Local Part
For the local part, we prove that
@@
Assume that the above claim does not hold. It indicates that we can find a sequence of mixing measures $G_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(c_{n,j^{\prime}})\delta_{{\bm{B}}_{n,j^{\prime}}{\bm{A}}_{n,j^{\prime}}}$ in $\bar{\mathcal{G}}_{L^{\prime}}(\Theta)$ such that
@@
Since $\mathcal{D}_{2n}\to 0$, we have $\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\to\exp(c_{*,j})$, ${\bm{Z}}_{n,i}\to{\bm{Z}}_{j}^{*}$ for any $i\in\mathcal{V}_{j}$ and $j\in[L]$. Throughout this proof, we assume without loss of generality that $M^{0}_{K,j}=I_{\bar{d}}$ with a note that our techniques can be extended to the general setting of that matrix. Now, the proof of the local part is divided into three steps as follows:
@@
Then, we can decompose the function $Q_{n}(\mathbb{X})$ as follows:
@@
We denote $\bar{U}(\mathbb{X};{\bm{Z}}):=\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}})\mathbb{X})$ and $\bar{V}(\mathbb{X};{\bm{Z}})=({\bm{M}}^{0}_{V}+{\bm{Z}})\mathbb{X}$, and $F(\mathbb{X};{\bm{Z}})=\bar{U}(\mathbb{X};{\bm{Z}})\bar{V}(\mathbb{X};{\bm{Z}})$. Based on the number of elements in each Voronoi cells, we decompose the function $\bar{A}_{n}(\mathbb{X})$ as follows:
@@
An application of the first-order Taylor expansion leads to
@@
for any $i$ and $j$ such that $i\in\mathcal{V}_{j}$ and $|\mathcal{V}_{j}|=1$. Here, the functions $\bar{R}_{ij,1}(\mathbb{X})$ and $\bar{R}_{ij,2}(\mathbb{X})$ denote the Taylor remainders. Collecting the above results leads to
@@
where the function $\bar{R}_{n,1}(\mathbb{X})$ satisfies that $\bar{R}_{n,1}(\mathbb{X})/\mathcal{D}_{2n}\to 0$. It is due to the uniform Lipschitz property of the function $F$. In the above display, the formulations of the coefficients $\bar{M}_{n,j,\alpha}$ are given by:
@@
Moving to the function $\bar{A}_{n,2}(\mathbb{X})$, an application of the Taylor expansion up to the second order leads to
@@
$\bar{A}_{n,1}(\mathbb{X})=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{V}+{\bm{Z}}_{j}^{*})\mathbb{X}+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}\Big{]}$
$\hskip 426.79134pt+\bar{R}_{n,1}(\mathbb{X}),$
$\bar{A}_{n,2}(\mathbb{X})=\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{V}+{\bm{Z}}_{j}^{*})\mathbb{X}+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}$
@@
$Q_{n}(\mathbb{X})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\exp(\mathbb{X}^{\top}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X})\Big{[}\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(u_{1})}\mathbb{X}^{(v_{1})}({\bm{M}}^{0}_{Q}+{\bm{Z}}_{j}^{*})\mathbb{X}+\sum_{u_{1},v_{1}=1}^{\bar{d}}\bar{M}_{n,j,e_{u_{1}v_{1}}}\mathbb{X}^{(v_{1})}e_{u_{1}}$
@@
Now, for any indices $j\in[L]$ such that $|\mathcal{V}_{j}|=1$, the limits $\bar{M}_{n,j,e_{uv}}/\mathcal{D}_{2n}\to 0$ lead to $\frac{\sum_{i\in\mathcal{V}_{j}}\exp(c_{n,i})\|\Delta{\bm{Z}}_{n,ij}\|_{1}}{\mathcal{D}_{2n}}\to 0$. That result directly implies that
@@
By putting the results in equations (26), (27), and (28) together, we arrive at $1=\frac{\mathcal{D}_{2n}}{\mathcal{D}_{2n}}\to 0$ as $n\to\infty$, which is a contradiction. As a consequence, at least one of the coefficients of the linear independent functions in $Q_{n}(\mathbb{X})/\mathcal{D}_{2n}$ does not go to 0 as $n\to\infty$.
Step 3 - Application of the Fatou’s lemma.
We denote $\bar{m}_{n}$ as the maximum of the absolute values of the coefficients of the linear independent functions in $Q_{n}(\mathbb{X})/\mathcal{D}_{2n}$. As at least one of these coefficients does not go to 0, it indicates that $1/\bar{m}_{n}\not\to\infty$ as $n\to\infty$. Since $\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/\mathcal{D}_{2n}\to 0$ as $n\to\infty$, we obtain $\|f_{G_{n}}-f_{G_{*}}\|_{L^{2}(\mu)}/(\bar{m}_{n}\mathcal{D}_{2n})\to 0$. An application of the Fatou’s lemma leads to:
$0=\lim_{n\to\infty}\dfrac{\|f_{G_{n}}-f_{\bar{G}_{*}}\|_{L^{2}(\mu)}}{\bar{m}_{n}\mathcal{D}_{2n}}\geq\int\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}(\mathbb{X})-f_{\bar{G}_{*}}(\mathbb{X})\right|}{\bar{m}_{n}\mathcal{D}_{2n}}d\mu(\mathbb{X})\geq 0$